OpenStudy (anonymous):

The amount of a chemical solution is measured to be 2 liters. What is the percent error of the measurement? I know the answer is 25% but I dont know how The solution is this. greatest possible error = 0.5(1) = 0.5 liter percent error = 0.5/2 = 0.25 Can anyone please explain?

OpenStudy (anonymous):

Question is missing something. There is only 1 variable. What are you trying to solve for? There must be some sort of other variable.

OpenStudy (anonymous):

that is the entire question. it is a quiz question but I dont know how they got the answer.

OpenStudy (anonymous):

greatest possible error = 0.5(1) = 0.5 liter percent error = 0.5/2 = 0.25 this is the solution